(6 marks)

Following are the transactions in Paris city last year:                                   

Item                                                                 AED

Taxes                                                              200,000

Investments                                                     300,000

Government Expenditure                                150,000

Exports                                                           250,000

Savings                                                            500,000

1.     Calculate government balance and explain the result.   (2 marks)

                                                                        

  1.  Calculate private sector balance.       (2 marks)                                                                                     

Calculate net exports and explain the result.   


1
Expert's answer
2021-01-19T07:30:41-0500

(a) government balance=200000-300000-150000=-250000

The city is in budget deficit. This is because the expenses exceed revenues.

(b)private sector balance=500000-300000=AED 200000

(c) net exports=250000-500000= AED-2500 000

The city is in a trade deficit .This is because it imports more than it exports